### v1

@@ 4.3 Hand Shape Reconstruction
Our framework uses the MANO model [MANO:SIGGRAPHASIA:2017] to represent human hands. This parametric model facilitates the detailed modeling of the human hand mesh using shape and pose parameters: the shape parameters $\beta\in\mathbb{R}^{10}$ capture individual hand identity, reflecting variations across different subjects, while the pose parameters $\theta\in\mathbb{R}^{51}$ describe the dynamic positions and orientations of the hand. We employ the method described in [19], a PyTorch-compatible differentiable implementation of MANO, enabling efficient optimization of hand poses and shapes. Prior to hand pose optimization, we undertake a pre-calibration process for the MANO shape parameters for each subject in our dataset.
@@ 4.4 Hand Pose Estimation
To overcome these challenges and improve the accuracy of our 3D keypoints, we adopt a RANSAC [FischlerBolles1981] filtering approach. For a hand joint, let us assume that it is detected by MediaPipe on $C_{\text{valid}}$ camera views. Then we can compute a set of candidate 3D keypoints, $\mathcal{X}_{i}=\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{valid}}}$, $N_{\text{valid}}=C_{\text{valid}}(C_{\text{valid}}-1)/2$ by triangulating pairs of valid camera views. A projection loss function is defined for a 3D keypoint:
@@ 5.1 Hand Pose Estimation
First, our dataset can be used to evaluate hand pose estimation. We provide annotations for 2D positions and 3D positions for the 21 hand joints of human hands in the dataset. In this experiment, we have evaluated two recent hand pose estimation methods: A2J-Transformer [21] and HaMeR [36], where ground truth bounding boxes of hands are used as input for hand pose estimation. The evaluation results are presented in Table 3. We used the PCK (Percentage of Correct Keypoints) metric for 2D hand pose estimation and the MPJPE (Mean Per Joint Position Error) metric for 3D hand pose estimation. A2J-Transformer extends A2J [46], a depth-based hand pose estimator, to use RGB images as input, and also uses a transformer architecture to encode non-local information. The A2J-Transformer model was trained on the InterHand2.6M dataset [32]. HaMeR [36] uses a large-scale ViT backbone [12] followed by a transformer decoder to regress the parameters of the hand. The HaMeR model was trained on 2.7M images collected from 10 existing hand pose datasets. We can see from Table 3, the performance of HaMeR is signficantly better than A2J-Transformer, largely due to using the ViT backbone and large-scale training images. The MPJPE of HaMeR is 29.1mm, which is much larger than the errors reported in other datasets in  [36]. This indicates that our dataset poses certain challenges for hand pose estimation, especially occlusions between hands and objects.

### v2

@@ 4.3 Hand Shape Reconstruction
Our framework uses the MANO model [38] to represent human hands. This parametric model facilitates the detailed modeling of the human hand mesh using shape and pose parameters: the shape parameters $\beta\in\mathbb{R}^{10}$ capture individual hand identity, reflecting variations across different subjects, while the pose parameters $\theta\in\mathbb{R}^{51}$ describe the dynamic positions and orientations of the hand. We employ the method described in [20], a PyTorch-compatible differentiable implementation of MANO, enabling efficient optimization of hand poses and shapes. Prior to hand pose optimization, we undertake a pre-calibration process for the MANO shape parameters for each subject in our dataset.
@@ 4.4 Hand Pose Estimation
To overcome these challenges and improve the accuracy of our 3D keypoints, we adopt a RANSAC [14] filtering approach. For a hand joint, let us assume that it is detected by MediaPipe on $C_{\text{valid}}$ camera views. Then we can compute a set of candidate 3D keypoints, $\mathcal{X}_{i}=\{\mathbf{x}_{i}\}_{i=1}^{N_{\text{valid}}}$, $N_{\text{valid}}=C_{\text{valid}}(C_{\text{valid}}-1)/2$ by triangulating pairs of valid camera views. A projection loss function is defined for a 3D keypoint:
@@ 4.5 Joint Hand-Object Pose Optimization
4.6 HoloLens Camera Pose Estimation
During the data collection process, we observed that head poses often experience jitter and do not move at a uniform speed. This makes it challenging for HoloLens to calculate accurate poses in real time, necessitating the optimization of the head poses published by HoloLens.
At each time step $t$, we have the HoloLens camera pose $\mathbf{T}_{t}^{H}$ in the world frame published from the device. To refine this camera pose, we obtain the optimized object poses $\mathcal{P}_{t}^{O}$ as described in Sec. 4.5. We treat the group of objects as a single merged object with pose $\mathbf{T}_{t}^{O}$ in the world frame. We can transform this object pose to the HoloLens camera frame denoted as $\mathbf{T}_{t}^{OH}$, assuming that the HoloLens frames and RealSense frames are well synchronized. Subsequently, we use FoundationPose [46] to optimize the object pose in the HoloLens camera frame given the initial object pose $\mathbf{T}_{t}^{OH}$, resulting in a refined HoloLens camera pose $\mathbf{T^{*}}_{t}^{H}$. Fig. 6 shows the comparison between the published pose and the optimized pose by projecting objects to the HoloLens camera frame, where the optimized poses demonstrate reduced jitter and more accurate object alignment.
Figure 6: Comparison between the published and refined HoloLens poses
@@ 5.1 Hand Pose Estimation
First, our dataset can be used to evaluate hand pose estimation. We provide annotations for 2D positions and 3D positions for the 21 hand joints of human hands in the dataset. In this experiment, we have evaluated two recent hand pose estimation methods: A2J-Transformer [22] and HaMeR [37], where ground truth bounding boxes of hands are used as input for hand pose estimation. The evaluation results are presented in Table 3. We used the PCK (Percentage of Correct Keypoints) metric for 2D hand pose estimation and the MPJPE (Mean Per Joint Position Error) metric for 3D hand pose estimation. A2J-Transformer extends A2J [48], a depth-based hand pose estimator, to use RGB images as input, and also uses a transformer architecture to encode non-local information. The A2J-Transformer model was trained on the InterHand2.6M dataset [33]. HaMeR [37] uses a large-scale ViT backbone [12] followed by a transformer decoder to regress the parameters of the hand. The HaMeR model was trained on 2.7M images collected from 10 existing hand pose datasets. We can see from Table 3, the performance of HaMeR is signficantly better than A2J-Transformer, largely due to using the ViT backbone and large-scale training images. The MPJPE of HaMeR is 28.9mm, which is much larger than the errors reported in other datasets in  [37]. This indicates that our dataset poses certain challenges for hand pose estimation, especially occlusions between hands and objects.